A 5.44 g
mixture contains both lithium fluoride, LiF,
and potassium fluoride, KF.
If the mixture contains 3.48 g
fluorine, what is the mass of the KF
in the mixture?
1
Expert's answer
2020-10-06T09:36:03-0400

system of equations:

58*x+26*y=5,44g

(x+y)*19=3, 48

where is x- chemical amount of KF and y is chemical amount of LiF

x=0,0211

m(KF)=0,0211*58=1,23g